OpenStudy (anonymous):

If f(x) = x2 - 1, find f(3). 5 2 8 -2 @eSpeX

OpenStudy (anonymous):

@jim_thompson5910

OpenStudy (anonymous):

f(0)= 0^2 -1 f(1) = 1^2 -1 f(2) = 2^2-1 see the pattern?

OpenStudy (anonymous):

Kind of, yes.

OpenStudy (anonymous):

@mathslover ?

OpenStudy (anonymous):

so the f(3) would be ?

OpenStudy (anonymous):

ummm f(3)= 2^3 -1?

OpenStudy (anonymous):

f(3) = x2 - 1 then solve

jimthompson5910 (jim_thompson5910):

All you're doing is replacing x with 3 f(x) = x^2 - 1 f(3) = 3^2 - 1 ... replace each copy of 'x' with 3 f(3) = ???

mathslover (mathslover):

is it : \(f(x) = x^2 - 1\) ?

OpenStudy (anonymous):

f(3) =8?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Thanks everyone!

mathslover (mathslover):

Jim, has got the point correct. You just have to put x = 3 for \(f(3) \) For example : \(f(x) = x^2 - 1\\ f(1) = 1^2 - 1 \\ f(2) = 2^2 - 1 \\ f(3) = ?\)